OpenStudy (maddiie):

Given the function defined by t(x)=x^3-3x^2+2x-3 t(-1)

OpenStudy (anonymous):

(-1)^3 - 3(-1)^2 + 2(1) -3 = -1-3+2-3 = -5

OpenStudy (maddiie):

Oh i get it so you insert the -1 in every spot that x goes

OpenStudy (maddiie):

Wouldn't it be -9 because you put a positive 1 with the 2 and it's actually negative

OpenStudy (maddiie):

\[(-1)^3-3(-1)^2+2(-1)-3= -1-3+-2-3=-9\]

OpenStudy (maddiie):

Or did i understand it wrong? Correct me if i am?

OpenStudy (anonymous):

Yes, you did it right. I'm sorry. I had made a mathematical error. the answer is -9

OpenStudy (maddiie):

Yay! So you taught me something. Awesome!! Thanks. I APPRECIATE IT

OpenStudy (anonymous):

Haha! No problem :)

OpenStudy (anonymous):

Thanks for correcting me!

OpenStudy (maddiie):

LOL. Well good night No problem i just actually want to learn not like other ppl that just get answers